# KiDS-Legacy: angular galaxy clustering from deep surveys with complex selection effects

### v1

@@ 5.2.1 Introduction to the method
With these assumptions and simplifications in mind, we find the optimal $NC_{\mathrm{KiDS}}$ by first choosing the same $NC_{\mathrm{KiDS}}$ and $NC_{\mathrm{rec}}$ values for data-driven OR and recovered mock OR, respectively. We vary the cluster number from low to high until it is high enough to pick up the LSS and the recovered 2PCF starts to be systematically lower than the unbiased case. On the other hand, if a certain choice of $NC$ gives an unbiased mock 2PCF, it is still possible that $NC_{\mathrm{KiDS}}$ is too low to recover the selection function, which results in a soft data-driven OR map that can be recovered with the same $NC_{\mathrm{rec}}=NC_{\mathrm{{KiDS}}}$. To test if this is the case, we can manually amplify the variance of the data-driven selection function that applies to the mock by:
@@
Note that $\mathcal{W}^{m}_{\mathrm{KiDS}}$ is $\mathcal{W}_{\mathrm{KiDS}}$ to the power of $m$. When the adjustment parameter $m>1$, the variance in $\delta_{\mathrm{OR,KiDS}}$ is enhanced. If the mock OR can still recover an unbiased 2PCF, then we can conclude that this setup is powerful enough to recover the realistic variable depth. In this test, we choose $m=1.5$ so that the bias in the 2PCF with URs is almost doubled.
@@ 5.2.2 Evaluating the performance of SOM+HC with data-driven systematics
Figure 13: Upper panel: part of the data-driven OR weight generated from the KiDS-Legacy map; middle panel: mock OR weight recovered from GLASS mock galaxy sample selected according to the data-driven OR; bottom panel: the relative difference between recovered OR weights and true OR weights. Both OR weights are generated with 600 hierarchical clusters and pixelised on a HEALPix map with Nside=2048.
Figure 13 shows an example of the true data-driven OR (the top panel), associated recovered OR (the bottom panel), and the relative difference between them. Both OR weights are generated with 600 hierarchical clusters and pixelized on a HEALPix map with Nside=2048. Visually they look very similar to each other, and the relative difference is well within ±$\sim\pm 20\%$.
Following Sec. 5.1, we quantitatively evaluate the effectiveness of SOM+HC by comparing the 2PCF from the depleted mock catalogue corrected by the recovered mock OR (the “Rec. OR $w(\theta)$”) and the “No selection” 2PCF (measured from a complete catalogue with the same number of galaxies, corrected by a uniform random). Since the “no selection” 2PCF is unbiased, it serves as the reference 2PCF. The consistency between these two 2PCFs indicates the consistency between the recovered mock OR and the data-driven OR (not necessarily between the recovered mock OR and the realistic selections). The angular separation is binned into 20 logarithmic bins from 2.5 arcmin to 250 arcmin. The covariance matrices are again calculated by the OneCovariance code with the same cosmology, redshift distribution, sky coverage and galaxy number density as the GLASS mock (see Figure 14).
@@
We can also evaluate the consistency at the level of parameter fits. For example, we set the matter density parameter $\Omega_{\mathrm{m}}$ and the galaxy bias $b$ as free parameters. First, we define a Gaussian likelihood:
@@
$\chi^{2}_{{\Omega}_{\mathrm{m}},b}=\begin{pmatrix}\Delta\bar{\Omega}_{\mathrm{m}}\\ \Delta\bar{b}\end{pmatrix}^{\rm{T}}\,\@tens{C}^{-1}_{\Omega_{\mathrm{m}},b}\,\begin{pmatrix}\Delta\bar{\Omega}_{\mathrm{m}}\\ \Delta\bar{b}\end{pmatrix}\,,$ (23)
where ${C}$ is the covariance matrix of the inferred parameters derived from the MCMC chains. The PTE value is then derived from the $\chi^{2}$ value given the degrees of freedom of 2.
@@ 5.3 Results
Figure 14: 2PCFs measured for the data-driven systematics test with the same $NC_{\mathrm{KiDS}}$ (the number of hierarchical clusters for the data-driven OR) and $NC_{\mathrm{rec}}$ (for the recovered mock OR). Each top panel corresponds to a setup. Nside is fixed at 2048 for both the data-driven OR and the recovered OR. The last panel is an enhanced data-driven OR with $m=1.5$ according to Eq. (21). The bottom panels show the relative difference between “recovered OR” and “no selection” 2PCFs. The data points are the mean 2PCF calculated from 40 GLASS realisations, and the error bars are calculated as the square root of the diagonal terms of the theoretical covariance matrix. The shaded regions are angular scales corresponding to a physical scale $r<8h^{-1}\mathrm{Mpc}$ at the mean redshift of the galaxy sample.
@@
Table 3: Summary statistics of data-driven systematics test. Four statistics are used to evaluate the consistency between 2PCFs and “No selection” 2PCFs. They are 1) $\chi^{2}_{\mathrm{d}}$ (defined in Eq. (18)) describes the difference between 2PCF data vectors; 2) probability-to-exceed (PTE) indicates the probability for a random data vector having higher $\chi^{2}_{\mathrm{d}}$ value; 3) $\Delta\Omega_{\mathrm{m}}$ and $\Delta b$ describes the individual parameter bias in the unit of constraining error; 4) $\chi^{2}_{\Delta\Omega_{\mathrm{m}},\,\Delta b}$ describes the overall constraining bias on the parameter space. The first 5 rows evaluate the difference between “Recovered OR” and “No selection” with the same $NC$ setup for the data-driven OR and recovered mock OR, and the following 3 rows evaluate the difference for fixed $NC_{\mathrm{KiDS}}$ test, and the last two rows evaluate the difference between “True OR”, “UR” cases and “No selection” cases, respectively;
@@
Figure 15: $1-\sigma$ confidence contours of the parameter posterior shift with respect to the best-fit values constrained from the “No selection” 2PCF. The pink arrow indicates the direction of the “UR” contour with the best-fit values of $\Delta\Omega_{\rm m}=-0.03$, $\Delta b=1.33$, which is well outside the dynamic range of the plot.
@@
In the upper panels of Figure 14, we present the 2PCFs measured from the data-driven systematics test. Each panel shows a setup specified by the number of clusters for data-driven OR and recovered mock OR. The last panel corresponds to an enhanced data-driven OR by setting $m=1.5$ in Eq (21). The colour scheme and meaning of data points are the same as introduced in Table 2. The data points are the mean 2PCF computed from 40 GLASS realisations and the error bars are the square root of the diagonal terms of the theoretical covariance matrix given by the OneCovariance code. The lower panels show the fractional bias of the 2PCFs with respect to the “no selection” 2PCF. The $1-\sigma$ confidence contours of the parameter posterior shift with respect to the best-fit values constrained from the “No selection” 2PCF are shown in Figure 15. Note that the “UR” 2PCF gives such biased constraints that the contour is well outside the plot area.
@@
Table 2 summarises the statistics defined in Subsection 5.2.2. The uniform random gives extremely biased 2PCF, highlighting the necessity to mitigate the variable depth. The “true OR” case agrees well with the “no selection” case as expected, and this means that using the true OR in Eq. 7 can indeed correct the variable depth for the 2PCF. From $NC=200$ to $NC=800$ the SOM+HC changes from undercorrecting the 2PCF to slightly overcorrecting it. The measured 2PCF and constrained cosmological parameters are in best agreement with the “no selection” case at $NC=600$. When we use an enhanced data-driven selection on the mock catalogue by setting $m=1.5$, $\chi_{\mathrm{d}}$ for UR increases from 8 47584758\,4758 475 (when $m=1$) to 20 5782057820\,57820 578, meaning that the bias in UR 2PCF is more than doubled when $m=1.5$. The SOM+HC method with $NC=600$ can still correct the bias in the 2PCF. Therefore, we conclude that SOM+HC with $NC\sim 600$ is the optimal choice for the KiDS-Legacy galaxy clustering 2PCF measurement.
@@
Now we fix $NC_{\mathrm{KiDS}}=600$ so that the data-driven selection reflects the realistic variable depth, then we vary $NC_{\mathrm{rec}}$ from 200 to 800. The evaluation statistics are summarised from the 6th row to the 8th row of Table 3, plus the third row with $NC_{\mathrm{rec}}=600$. Again we see that the recovered 2PCF goes from biased high (when $NC_{\mathrm{OR}}=200$) to well corrected. We also note that the recovery biases are generally low with $NC_{\mathrm{rec}}$ varying from 400 to 800, for which $\chi^{2}_{\mathrm{d}}$ values are at a level of $\sim 0.5\sigma$, with a minimum value of 0.39 when $NC_{\mathrm{rec}}=600$. The goodness of recovery is also robust to the choice of $NC_{\mathrm{rec}}$ from 400 to 800. In conclusion, we will choose $NC_{\mathrm{rec}}=600$ for the galaxy clustering measurement with KiDS-Legacy data.
@@ 6.1 Application of SOM+HC on the real data
Figure 16: Self-organising maps trained on the KiDS-Legacy catalogue, with the dimension $30\times 30$. The first five panels are SOMs coloured by the average systematics values in each cell. The last panel (bottom right) is the SOM coloured by the galaxy number contrast of each hierarchical cluster. The black lines are the boundaries of each HC. Note that we use a toroidal topology for the SOM, so the left and right edges and the top and bottom edges are adjacent.
@@
In this section, we apply the SOM+HC method to the real KiDS-Legacy galaxy catalogue. The systematics used to train the SOM are described in Sec. 3. The SOM+HC setup is the same as the data-driven systematics test: the SOM dimension is 30, with hexagonal cells (so each cell has six adjacent neighbourhoods), toroidal topology (so the left and right edges are adjacent, and the top and bottom edges are adjacent). Training lasts for 10 epochs, with an initial width of the Gaussian neighbourhood function $\sigma=30/2=15$, decreasing linearly to $\sigma=1$ in the last epoch. The initial learning rate is 0.1 and decreases linearly to 0.01. The trained SOM cells are grouped into 600 clusters, from which the OR weight is generated on a HEALPix map with Nside=2048.
Figure 16 shows the results of the post-trained SOM+HC. The first five panels show the SOM colour-coded by the average systematics values of each cell. Note that the SOM inherits the grained patterns of the mean Level and GAIA_nstar spatial distributions (see Figure 3). The last panel shows the galaxy number contrast of the 600 hierarchical clusters (denoted by cells with black borders). Due to the toroidal topology, clusters near the edge may actually cross the edge and continue on the other side. One can see the correlation between galaxy number contrast and each systematics from the SOM maps. For example, the galaxy number contrast is strongly anticorrelated with PSF_size. This is expected, as a larger PSF size indicates poorer atmospheric seeing, making galaxies harder to detect under these conditions. A quantitative correlation can be evaluated using Spearman correlation coefficients. Figure 17 shows the correlation coefficient matrix. The bottom row shows the correlation between systematics and galaxy number contrast, and it is clear that PSF size has a strong anti-correlation with galaxy number contrast. On the other hand, the PSF shape only has a weak correlation with the contrast in galaxy numbers. We also find that Level and GAIA_nstar are correlated with extinction.
@@ 6.2 Blinded measurement of 2PCF
The 2PCF is measured the same way as those from Section 5. The “UR” 2PCF is measured by using the coverage map as the random term $N^{a}_{R,p}$ in Eq (9) (for the similar definition of the $RR$ and $DR$ terms), and the “OR” 2PCF uses the OR weight from SOM+HC. Both 2PCFs are then blinded using the method given in Muir et al. (2020). The blinding method includes the following steps:
@@
    Select a reference cosmology and calculate the associated theoretical 2PCF. In our case, we use the fiducial cosmology defined in Section 1.
@@
Both the UR and OR 2PCFs are shifted by the same amount to make a meaningful comparison. We then run an MCMC on the blinded 2PCFs to sample the posterior of $\Omega_{\mathrm{m}}$, $b$}. To prevent accidental unblinding, we never save the blinded data vector, but directly pass it to the MCMC code. The covariance matrix is computed from the OneCovariance code. Since we do not know the true parameter values a priori, we use an iterative fitting procedure: first, we compute the covariance matrix with the fiducial cosmology and run MCMCs with it, then we compute the best-fit parameters from the posterior and update the covariance matrices for OR and UR respectively. The posterior is then sampled with the updated covariance matrices. The MCMCs are sampled with the same modelling code when we blind the data vector. Note that we only constrain the linear model with data points at $\theta>8h^{-1}\mathrm{Mpc}/D_{A}(\bar{z})$, where $D_{A}(\bar{z})$ is the angular diameter distance at the mean KiDS-Legacy redshift $\bar{z}=0.7$.
@@
Data-driven systematics tests (see Section 5) suggest that $NC=600$ is the optimal choice. In this section, we use this as the fiducial setting, but also try $NC=\{200,400,800\}$ to test the robustness of the $NC$ choice. All 2PCFs are shifted with the same blinding shift. The posteriors are also fitted from these measurements.
@@
The blinded 2PCFs are shown in Figure 18 with error bars derived from the updated covariance matrices. The black dashed line shows the theoretical 2PCF calculated from the best-fit parameters from the blinded 2PCF measurement with $NC=600$. The reduced $\chi^{2}$ value between the OR ($NC=600$) 2PCF and the best-fit 2PCF is 1.28, corresponding to a PTE of 0.25, indicating a good fit between the model and the blinded data. Figure 19 shows the posteriors sampled from the MCMC, shifted with respect to the best-fit value of the $NC=600$ case. The left panel shows the contours of shifted posteriors constrained from 2PCF corrected by OR with different $NC$, while the right panel shows contours corresponding to 2PCF corrected by OR ($NC=600$) and UR, respectively. On linear scales, the OR and UR 2PCFs differ at a level of $70\sigma$ and the galaxy bias parameter differs at a level of $40\sigma$. Interestingly, the matter density $\Omega_{\mathrm{m}}$ does not change significantly with the correction. This is because the selection does not cause significant bias in $\Omega_{\mathrm{m}}$ (note that $\Omega_{\mathrm{m}}$ is not biased even for the UR 2PCF), but this might be not the case for other surveys.
@@
The 2PCF measurements with different $NC$ suggest that the OR 2PCFs decrease with $NC$, i.e. the corrections increase with $NC$. If we treat $NC=600$ as the fiducial choice, the corresponding $\chi_{\mathrm{d}}$ values are 5.75, 0.27, 0.07 with $NC=200,400,800$, so the change is significant from $NC=200$ to $NC=400$, while moderate from $NC=400$ to $NC=800$. This is consistent with the data-driven validation test. Therefore, we conclude that the OR correction is quite robust around $NC=600$.
@@
Figure 18: The blinded 2PCFs measured from the KiDS-Legacy sample. Blue and pink dots are the measurements corrected by the recovered organised random, and by the uniform random, respectively. The shaded regions are angular scales corresponding to physical scales smaller than $8h^{-1}\mathrm{Mpc}$ estimated at the mean redshift. The error bars are the standard deviations derived from the covariance matrix computed by the OneCovariance code. The black dashed curve shows the best-fit 2PCF from the MCMC.
@@
Figure 19: Contours of 68.3% and 95.4% credible levels of the parameter posterior shift with respect to the best-fit values constrained from $NC=600$ case. Left panel: contours from the OR 2PCFs with different $NC$ choices. The fiducial choice $NC=600$ is shown as green filled contour; Right panel: contour of OR 2PCF ($NC=600$, green contour) and UR 2PCF (pink contour).

### v2

@@ 5.2.1 Introduction to the method
With these assumptions and simplifications in mind, we find the optimal $NC_{\mathrm{KiDS}}$ by first choosing the same $NC_{\mathrm{KiDS}}$ and $NC_{\mathrm{rec}}$ values for data-driven OR and recovered mock OR, respectively. We vary the cluster number from low to high until it is high enough to pick up the LSS and the recovered 2PCF starts to be systematically lower than the unbiased case. On the other hand, if a certain choice of $NC$ gives an unbiased mock 2PCF, it is still possible that $NC_{\mathrm{KiDS}}$ is too low to recover the selection function, which results in a soft data-driven OR map that can be recovered with the same $NC_{\mathrm{rec}}=NC_{\mathrm{{KiDS}}}$. To test if this is the case, we can manually amplify the variance of the data-driven selection function that applies to the mock by
@@
We note that $\mathcal{W}^{m}_{\mathrm{KiDS}}$ is $\mathcal{W}_{\mathrm{KiDS}}$ to the power of $m$. When the adjustment parameter $m>1$, the variance in $\delta_{\mathrm{OR,KiDS}}$ is enhanced. If the mock OR can still recover an unbiased 2PCF, then we can conclude that this set-up is powerful enough to recover the realistic variable depth. In this test, we choose $m=1.5$ so that the bias in the 2PCF with URs is almost doubled.
@@ 5.2.2 Evaluating the performance of SOM+HC with data-driven systematics
Figure 13: Upper panel: Part of the data-driven OR weight generated from the KiDS-Legacy map; middle panel: Mock OR weight recovered from GLASS mock galaxy sample selected according to the data-driven OR; bottom panel: Relative difference between recovered OR weights and true OR weights. Both OR weights are generated with 600 hierarchical clusters and pixelised on a HEALPix map with Nside=2048.
Figure 13 shows an example of the true data-driven OR (the top panel), associated recovered OR (the bottom panel), and the relative difference between them. Both OR weights are generated with 600 hierarchical clusters and pixelised on a HEALPix map with Nside=2048. Visually they look very similar to each other, and the relative difference is well within ±$\sim\pm 20\%$.
Following Sec. 5.1, we quantitatively evaluate the effectiveness of SOM+HC by comparing the 2PCF from the depleted mock catalogue corrected by the recovered mock OR (the “Rec. OR $w(\theta)$”) and the “No selection” 2PCF (measured from a complete catalogue with the same number of galaxies, corrected by a uniform random). Since the “No selection” 2PCF is unbiased, it serves as the reference 2PCF. The consistency between these two 2PCFs indicates the consistency between the recovered mock OR and the data-driven OR (not necessarily between the recovered mock OR and the realistic selections). The angular separation is binned into 20 logarithmic bins from 2.5 arcmin to 250 arcmin. The covariance matrices are again calculated by the OneCovariance code with the same cosmology, redshift distribution, sky coverage and galaxy number density as the GLASS mock (see Fig. 14).
@@
We can also evaluate the consistency at the level of parameter fits. For example, we set the matter density parameter $\Omega_{\mathrm{m}}$ and the galaxy bias $b$ as free parameters. First, we define a Gaussian likelihood,
@@
$\chi^{2}_{{\Omega}_{\mathrm{m}},b}=\begin{pmatrix}\Delta\bar{\Omega}_{\mathrm{m}}\\ \Delta\bar{b}\end{pmatrix}^{\rm{T}}\,\@tens{C}^{-1}_{\Omega_{\mathrm{m}},b}\,\begin{pmatrix}\Delta\bar{\Omega}_{\mathrm{m}}\\ \Delta\bar{b}\end{pmatrix}\,.$ (23)
Here ${C}$ is the covariance matrix of the inferred parameters derived from the MCMC chains. The PTE value is then derived from the $\chi^{2}$ value given the degrees of freedom of 2.
@@ 5.3 Results
Figure 14: 2PCFs measured for the data-driven systematics test with the same choices of $NC_{\mathrm{KiDS}}$ (the number of hierarchical clusters for the data-driven OR) and $NC_{\mathrm{rec}}$ (for the recovered mock OR) Each column of panels corresponds to one set-up. Nside is fixed at 2048 for both the data-driven OR and the recovered OR. The right-most column is an enhanced data-driven OR with $m=1.5$ according to Eq. (21). The top panels show the 2PCF data points calculated as the mean values from 40 GLASS realisations, and the error bars are calculated as the square root of the diagonal terms of the theoretical covariance matrix. The middle panels show the biases in 2PCFs with respect to the No selection 2PCFs and the bottom panels show the biases relative to the errors in the 2PCFs (The UR case is well beyond the range). The shaded regions are angular scales corresponding to a physical scale $r<8h^{-1}\mathrm{Mpc}$ at the mean redshift of the galaxy sample.
Table 3: Summary statistics of data-driven systematics test.
@@
 [^18] Four statistics are used to evaluate the consistency between 2PCFs and “No selection” 2PCFs. They are 1) $\chi^{2}_{\mathrm{d}}$ (defined in Eq. (18)) describes the difference between 2PCF data vectors; 2) probability-to-exceed (PTE) indicates the probability for a random data vector having higher $\chi^{2}_{\mathrm{d}}$ value; 3) $\Delta\Omega_{\mathrm{m}}$ and $\Delta b$ describes the individual parameter bias in the unit of constraining error; 4) $\chi^{2}_{\Delta\Omega_{\mathrm{m}},\,\Delta b}$ describes the overall constraining bias on the parameter space. The first five rows evaluate the difference between “Recovered OR” and “No selection” with the same $NC$ set-up for the data-driven OR and recovered mock OR, and the following three rows evaluate the difference for fixed $NC_{\mathrm{KiDS}}$ test, and the last two rows evaluate the difference between “True OR”, “UR” cases, and “No selection” cases, respectively;
@@
Figure 15: $1-\sigma$ confidence contours of the parameter posterior shift with respect to the best-fit values constrained from the No selection 2PCF. The pink arrow indicates the direction of the UR contour with the best-fit values of $\Delta\Omega_{\rm m}=-0.03$, $\Delta b=1.33$, which is well outside the dynamic range of the plot.
@@
In the upper panels of Fig. 14, we present the 2PCFs measured from the data-driven systematics test. Each column of panels shows a set-up specified by the number of clusters for data-driven OR and recovered mock OR. The right-most column corresponds to an enhanced data-driven OR by setting $m=1.5$ in Eq (21). The colour scheme and meaning of data points are the same as introduced in Table 16. The data points are the mean 2PCF computed from 40 GLASS realisations and the error bars are the square root of the diagonal terms of the theoretical covariance matrix given by the OneCovariance code. The middle panels show the fractional bias of the 2PCFs with respect to the “No selection” 2PCF and the bottom panels show the bias of the 2PCFs with respect to the error. The $1\sigma$ confidence contours of the parameter posterior shift with respect to the best-fit values constrained from the “No selection” 2PCF are shown in Fig. 15. We note that the “UR” 2PCF gives such biased constraints that the contour is well outside the plot area.
@@
Table 16 summarises the statistics defined in Sect. 5.2.2. The uniform random gives extremely biased 2PCF, highlighting the necessity to mitigate the variable depth. The “True OR” case agrees well with the “No selection” case as expected, and this means that using the true OR in Eq. 7 can indeed correct the variable depth for the 2PCF. From $NC=200$ to $NC=800$ the SOM+HC changes from under-correcting the 2PCF to slightly over-correcting it. The measured 2PCF and constrained cosmological parameters are in best agreement with the “No selection” case at $NC=600$. When we use an enhanced data-driven selection on the mock catalogue by setting $m=1.5$, $\chi_{\mathrm{d}}$ for UR increases from 8 47584758\,4758 475 (when $m=1$) to 20 5782057820\,57820 578, meaning that the bias in UR 2PCF is more than doubled when $m=1.5$. The SOM+HC method with $NC=600$ can still correct the bias in the 2PCF. Therefore, we conclude that SOM+HC with $NC\sim 600$ is the optimal choice for the KiDS-Legacy galaxy clustering 2PCF measurement.
@@
Now we fix $NC_{\mathrm{KiDS}}=600$ so that the data-driven selection reflects the realistic variable depth, then we vary $NC_{\mathrm{rec}}$ from 200 to 800. The evaluation statistics are summarised from the 6th row to the 8th row of Table 18, plus the third row with $NC_{\mathrm{rec}}=600$. Again we see that the recovered 2PCF goes from biased high (when $NC_{\mathrm{OR}}=200$) to well corrected. We also note that the recovery biases are generally low with $NC_{\mathrm{rec}}$ varying from 400 to 800, for which $\chi^{2}_{\mathrm{d}}$ values are at a level of $\sim 0.5\sigma$, with a minimum value of 0.39 when $NC_{\mathrm{rec}}=600$. The goodness of recovery is also robust to the choice of $NC_{\mathrm{rec}}$ from 400 to 800. In conclusion, we choose $NC_{\mathrm{rec}}=600$ for the galaxy clustering measurement with KiDS-Legacy data.
@@ 6.1 Application of SOM+HC on the real data
Figure 16: Self-organising maps trained on the KiDS-Legacy catalogue, with the dimension $30\times 30$. The first five panels are SOMs coloured by the average systematics values in each cell. The last panel (bottom right) is the SOM coloured by the galaxy number contrast of each hierarchical cluster. The black lines are the boundaries of each HC. We note that we use a toroidal topology for the SOM, so the left and right edges and the top and bottom edges are adjacent.
@@
In this section we apply the SOM+HC method to the real KiDS-Legacy galaxy catalogue. The systematics used to train the SOM are described in Sec. 3. The SOM+HC set-up is the same as the data-driven systematics test: the SOM dimension is 30, with hexagonal cells (so each cell has six adjacent neighbourhoods), toroidal topology (so the left and right edges are adjacent, and the top and bottom edges are adjacent). Training lasts for 10 epochs, with an initial width of the Gaussian neighbourhood function $\sigma=30/2=15$, decreasing linearly to $\sigma=1$ in the last epoch. The initial learning rate is 0.1 and decreases linearly to 0.01. The trained SOM cells are grouped into 600 clusters, from which the OR weight is generated on a HEALPix map with Nside=2048.
Figure 16 shows the results of the post-trained SOM+HC. The first five panels show the SOM colour-coded by the average systematics values of each cell. We note that the SOM inherits the grained patterns of the mean Level and GAIA_nstar spatial distributions (see Fig. 2). The last panel shows the galaxy number contrast of the 600 hierarchical clusters (denoted by cells with black borders). Due to the toroidal topology, clusters near the edge may actually cross the edge and continue on the other side. One can see the correlation between galaxy number contrast and each systematics from the SOM maps. For example, the galaxy number contrast is strongly anticorrelated with PSF_size. This is expected, as a larger PSF size indicates poorer atmospheric seeing, making galaxies harder to detect under these conditions. A quantitative correlation can be evaluated using Spearman correlation coefficients. Figure 17 shows the correlation coefficient matrix. The bottom row shows the correlation between systematics and galaxy number contrast, and it is clear that PSF size has a strong anti-correlation with galaxy number contrast, in agreement with the contrast-systematics relation shown as black curves in Fig. 2. On the other hand, the PSF shape only has a weak correlation with the contrast in galaxy numbers. We also find that Level and GAIA_nstar are correlated with extinction.
@@ 6.2 Blinded measurement of 2PCF
The 2PCF is measured the same way as those from Sect. 5. The “UR” 2PCF is measured by using the coverage map as the random term $N^{a}_{R,p}$ in Eq (9) (for the similar definition of the $RR$ and $DR$ terms), and the “OR” 2PCF uses the OR weight from SOM+HC. Both 2PCFs are then blinded using the method given in Muir et al. (2020). The blinding method includes the following steps:
@@
    Select a reference cosmology and calculate the associated theoretical 2PCF. In our case, we use the fiducial cosmology defined in Sect. 1.
@@
Both the UR and OR 2PCFs are shifted by the same amount to make a meaningful comparison. We then run an MCMC on the blinded 2PCFs to sample the posterior of $\Omega_{\mathrm{m}}$, $b$}. To prevent accidental unblinding, we never save the blinded data vector, but directly pass it to the MCMC code. The covariance matrix is computed from the OneCovariance code. Since we do not know the true parameter values a priori, we use an iterative fitting procedure: first, we compute the covariance matrix with the fiducial cosmology and run MCMCs with it, then we compute the best-fit parameters from the posterior and update the covariance matrices for OR and UR respectively. The posterior is then sampled with the updated covariance matrices. The MCMCs are sampled with the same modelling code when we blind the data vector. We note that we only constrain the linear model with data points at $\theta>8h^{-1}\mathrm{Mpc}/D_{A}(\bar{z})$, where $D_{A}(\bar{z})$ is the angular diameter distance at the mean KiDS-Legacy redshift $\bar{z}=0.7$. The redshift distribution is calibrated with a combination of SOM and the clustering-redshift method, which will also be presented in a companion paper.
@@
Data-driven systematics tests (see Sect. 5) suggest that $NC=600$ is the optimal choice. In this section, we use this as the fiducial setting, but also try $NC=\{200,400,800\}$ to test the robustness of the $NC$ choice. All 2PCFs are shifted with the same blinding shift. The posteriors are also fitted from these measurements.
@@
The blinded 2PCFs are shown in Fig. 18 with error bars derived from the updated covariance matrices. The black dashed line shows the theoretical 2PCF calculated from the best-fit parameters from the blinded 2PCF measurement with $NC=600$. The reduced $\chi^{2}$ value between the OR ($NC=600$) 2PCF and the best-fit 2PCF is 1.28, corresponding to a PTE of 0.25, indicating a good fit between the model and the blinded data. Figure 19 shows the posteriors sampled from the MCMC, shifted with respect to the best-fit value of the $NC=600$ case. The left panel shows the contours of shifted posteriors constrained from 2PCF corrected by OR with different $NC$, while the right panel shows contours corresponding to 2PCF corrected by OR ($NC=600$) and UR, respectively. On linear scales, the OR and UR 2PCFs differ at a level of $70\sigma$ and the galaxy bias parameter differs at a level of $40\sigma$. Interestingly, the matter density $\Omega_{\mathrm{m}}$ does not change significantly with the correction. This is because the selection does not cause significant bias in $\Omega_{\mathrm{m}}$ ($\Omega_{\mathrm{m}}$ is not biased even for the UR 2PCF), but this might be not the case for other surveys.
@@
The 2PCF measurements with different $NC$ suggest that the OR 2PCFs decrease with $NC$ (i.e. the corrections increase with $NC$). If we treat $NC=600$ as the fiducial choice, the corresponding $\chi_{\mathrm{d}}$ values are 5.75, 0.27, 0.07 with $NC=200,400,800$, so the change is significant from $NC=200$ to $NC=400$, while moderate from $NC=400$ to $NC=800$. This is consistent with the data-driven validation test. Therefore, we conclude that the OR correction is quite robust around $NC=600$.
@@
Figure 18: Blinded 2PCFs measured from the KiDS-Legacy sample. The blue and pink dots are the measurements corrected by the recovered organised random and by the uniform random, respectively. The shaded regions are angular scales corresponding to physical scales smaller than $8h^{-1}\mathrm{Mpc}$ estimated at the mean redshift. The error bars are the standard deviations derived from the covariance matrix computed by the OneCovariance code. The black dashed curve shows the best-fit 2PCF from the MCMC.
@@
Figure 19: Contours of 68.3% and 95.4% credible levels of the parameter posterior shift with respect to the best-fit values constrained from $NC=600$ case. Left panel: Contours from the OR 2PCFs with different $NC$ choices. The fiducial choice $NC=600$ is shown as green filled contours; Right panel: Contours of OR 2PCF ($NC=600$, green) and UR 2PCF (pink).